Find the derivative of y = arcsinh(1-x/1+x)derivatives

Grade:12

1 Answers

Askiitians Expert Bharath-IITD
14 years ago

Dear Satyam,

                    we know that d/dx(arcsinh x) = 1/√(x2+1)

So here in place of x we have (1-x/1+x)

And d/dx(1-x/1+x)= [ (1+x) d/dx(1-x) - (1-x) d/dx(1+x) ] / (1+x)2

                           = [ (1+x) *(-1) - (1-x) *1 ] / (1+x)2

                           = -2/ (1+x)2

now dy/dx = 1/ √((1-x/1+x)2+1) *  d/dx(1-x/1+x)

                =[ (1+x)/√2(x2+1) ] *  -2/ (1+x)2

                      = -√2/ √(x2+1)*(1+x)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  !!!

 


Regards,

Askiitians Experts

Adapa Bharath